OpenStudy (anonymous):

The last one for tonight Consider the functions f(x) = 3x2 and g(x) = 2(x + 9). Find f[g(-7)]. Please show all of your work for full credit.

OpenStudy (anonymous):

help!!! T_T

OpenStudy (anonymous):

plug -7 for the function of g(x) then the answer for that function goes into f(x) g(-7)=2(-7+9) f(g(x)) = 3( 2(-7+9))\[^{2}\]

OpenStudy (anonymous):

f(g(x)) = 3( 2(-7+9))^2

OpenStudy (anonymous):

is that all? :c

OpenStudy (anonymous):

yes just work out the equation 3( 2(-7+9))^2 add -7 and 9 then multiply the answer by 2 then square it then multiply it by 3 and thats the final answer

OpenStudy (anonymous):

thanks! C: